OpenStudy (anonymous):

Show that if f(x) is a differentiable function with f(x)<0 for all x E R and with a local maximum at x=c, then g(x)=[f(x)]^2 has a local minimum at x=c?

OpenStudy (zarkon):

if f(c) is a max then f'(c)=0 and the 1st d-test would give us the just to the left of c f' would be positive and just to the right of c f' would be negative g'(x)=2f(x)f'(x) g'(c)=2f(c)f'(c)=2f(c)*0=0 just to the left of c g' will be negative since f<0 and f'>0 just to the right of c g' will be positive since f<0 and f'<0 thus g(c) is a min

OpenStudy (anonymous):

how is f'(c)=0?

OpenStudy (anonymous):

Zarkon, would this also suffice? (im trying to get the hang of Analysis) f(c) being the max means: \[f(x)\leq f(c)<0 \] for all x in whatever interval we are talking about By squaring both sides we obtain: \[f(x) \leq f(c) 0 \Rightarrow 0 < (f(c))^2 \leq (f(x))^2\]

OpenStudy (anonymous):

oops, that should be f(c) < 0 on the left side.

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

cool. im gonna look at yours in more detail though.

OpenStudy (zarkon):

you would need to specify a neighborhood though (not for all x) this is a local max

OpenStudy (anonymous):

ah i gotcha. and i like your proof better, very elegant :)

OpenStudy (zarkon):

ok i see it ;)

OpenStudy (zarkon):

@ joe : have you taken real analysis yet?

OpenStudy (anonymous):

ive taken one semester in it, but tbh, i dont feel like i learned anything =/

OpenStudy (anonymous):

i read the book every now and then (when i can pry myself away from my Linear Algebra books lol)

OpenStudy (anonymous):

Thanks guys, but can you go through more word problems that I have posted. It will be great! =]

OpenStudy (zarkon):

it takes a while for it all to sink in. My first course in RA kicked my butt..after about the 5th class I was ok ;)

OpenStudy (anonymous):

Wow i dont even know if my school offers that many courses o.O It must be good though. I have a friend in graduate school that swears by Real Analysis lol

OpenStudy (zarkon):

it gets neat when you learn measure theory and functional analysis

OpenStudy (zarkon):

I had 1 class in RA as an undergrad and 5 as a grad

OpenStudy (anonymous):

oh, i never thought of looking in the grad catalog. Maybe after RA II i'll see what grad classes are available.

OpenStudy (zarkon):

you should ... it's good stuff

myininaya (myininaya):

i cant believe joe looked at this problem he doesnt like cal

OpenStudy (anonymous):

its not cal though! (not really) >.< and notice i didnt use Cal to solve it :P

myininaya (myininaya):

i guess lol

OpenStudy (anonymous):

i guess it is kinda a cal problem lol

OpenStudy (anonymous):

but i stick by my non-cal solution! rawr.

myininaya (myininaya):

i would had done it zarkon's way yours is pretty too though

OpenStudy (anonymous):

my way isnt as interesting though. I didnt use all the facts presented (which sorta makes me feel like its wrong)

OpenStudy (zarkon):

you don't need all the facts

OpenStudy (anonymous):

that makes me feel better lol

OpenStudy (zarkon):

your way works even if the function is not differentiable at c

OpenStudy (anonymous):

im not feeling too good >.< i think i just need some food. I'll see you guys later :)

OpenStudy (zarkon):

later

myininaya (myininaya):

poor joe